OpenStudy (anonymous):

y = 3x + 3 y = x − 1 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form.

TheSmartOne (thesmartone):

Any ideas?

TheSmartOne (thesmartone):

Do you know what the substitution method is?

TheSmartOne (thesmartone):

and not to forget... \(\bf\huge~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8

OpenStudy (anonymous):

just flip it to the other side y=3x+3 becomes y-3/3=x and y=x-1 becomes y+1=x

OpenStudy (anonymous):

thanksss

OpenStudy (texaschic101):

y = 3x + 3......so now we substitute 3x + 3 in for y in the other equation y = x - 1 3x + 3 = x - 1 --- subtract x from both sides 3x - x + 3 = x - x - 1 -- combine like terms 2x + 3 = -1 --- subtract 3 from both sides 2x + 3 - 3 = -1 - 3 -- combine like terms 2x = -4 -- divide both sides by 2 2/2x = -4/2 x = -2 now we sub in -2 for x in either of the original equations to find y y = 3x + 3 y = 3(-2) + 3 y = -6 + 3 y = -3 now we check.. y = x - 1 -3 = -2 -1 -3 = -3 (correct) so your solution is : (-2,-3)